How many milliliters of 0.100 M HClO₃ are required to neutralize 40.0 mL of 0.140 M KOH?
Answer #1

Balanced chemical equation is:

KOH + HClO3 ---> KClO3 + H2O

Here:

M(KOH)=0.14 M

M(HClO3)=0.1 M

V(KOH)=40.0 mL

According to balanced reaction:

1*number of mol of KOH =1*number of mol of HClO3

1*M(KOH)*V(KOH) =1*M(HClO3)*V(HClO3)

1*0.14 M *40.0 mL = 1*0.1M *V(HClO3)

V(HClO3) = 56.0 mL

Answer: 56.0 mL